Translate into algebraic equations :
1.The sum of three consecutive even numbers is 144; what are the numbers?
2.The price of one brand of athletic shoes is $5 more than another brand. If the prices of the two pairs of shoes add to $230, how much does each pair cost?
3. A man buys five DVDs for $66.34, including 7% sales tax. How much did each DVD cost?

Answer by rajagopalan(174)

Translate into algebraic equations :
1.The sum of three consecutive even numbers is 144; what are the numbers?
Let the mid number be n
the numbers are n-2, n and n+2
giving 3n=144
Therfore numbers are = (sum/3)-2, Sum/3 and sum/3+2
2.The price of one brand of athletic shoes is $5 more than another brand. If the prices of the two pairs of shoes add to $230, how much does each pair cost?
let x is cost of one brand
cost of other brand =x+5
x+x+5=sum=230
2x=sum-5
x=(sum-5)/2